The following information applies to the questions displayed below.]

Diego Company manufactures one product that is sold for $78 per unit in two geographic regions—the East and West regions. The following information pertains to the company’s first year of operations in which it produced 60,000 units and sold 57,000 units.

Variable costs per unit:
Manufacturing:
Direct materials $ 28
Direct labor $ 12
Variable manufacturing overhead $ 2
Variable selling and administrative $ 3
Fixed costs per year:
Fixed manufacturing overhead $ 1,260,000
Fixed selling and administrative expenses $ 654,000

The company sold 42,000 units in the East region and 15,000 units in the West region. It determined that $340,000 of its fixed selling and administrative expenses is traceable to the West region, $290,000 is traceable to the East region, and the remaining $24,000 is a common fixed cost. The company will continue to incur the total amount of its fixed manufacturing overhead costs as long as it continues to produce any amount of its only product.   

9.

Award: 0 out of 5.00 points

4. What is the company’s net operating income (loss) under variable costing?

8a. What is the company’s break-even point in unit sales?
8b. Is it above or below the actual sales volume?
Answer #1

Answers :

4) Variable costing net income = -33,000

8) A) BEP Units = 58,000 , B) Above the Actula Sales Volume

8A) Break Even Point in Units = fixed Cost / CM per unit

CM Per unit = Selling Price - Varible cost per unit = 78-45 =33

Fixed Cost = 1,260,000+654,000 =1,914,000

BEP in Units = 1,914,000/33 = 58,000 Units

Is it Above or BElow the Actula Sales : Above

Assuming the Firm Using the Variable Cositng Variable Costing Calculation of Unit cost Under Variable Costing Direct MaterialNote For Any Further Clarification Plases Use the Comment Box.Thank You Very much